OpenStudy (jazzyfa30):

The Jackson Band Boosters sell beverages for $1.75 and candy for $1.50 at home games. Their goal is to have total sales of $525 for each game. Write an equation in standard form that would model the different amounts if candy and drinks that could be sold to reach the goal. Use the variables x for beverages and y for candy

OpenStudy (jazzyfa30):

@3HUNNAREEK11

OpenStudy (chillout):

The total sales should be of $525. Therefore, the right side of the equation should be 525. Now for the left side: We can sell an amount 'a' of beverages and an amount 'b' of candies. Thus, we have\[1.75x + 1.5y = 525\]This is the standard form. Writing in terms of y, we have y = 350 - 7x/6.

OpenStudy (chillout):

Oops. It should read amount 'x' and amount 'y'

OpenStudy (jazzyfa30):

what???

OpenStudy (anonymous):

solve the inequality -y<-2

OpenStudy (jazzyfa30):

y>2

OpenStudy (anonymous):

solve the inequality -13>x-4

OpenStudy (jazzyfa30):

-9>x

OpenStudy (anonymous):

thanks

OpenStudy (chillout):

Well, I dunno how I can explaint it. But it does model the different values that y and x can assume. For example, if you sell 1 candy, you have 1.75x + 1.5 = 525, x = 523.5/1.75 Which is not a "true" [But it's still a solution] solution because it is not an integer.

OpenStudy (jazzyfa30):

ummmmm.....Okay I guess

OpenStudy (chillout):

If you set y=2 you'd have another solution for x. Do you want me to draw a graph?

OpenStudy (anonymous):

whats the soultion for 5 + y<-5

OpenStudy (jazzyfa30):

your welcome @Jazzy92159

OpenStudy (jazzyfa30):

y<-10

OpenStudy (jazzyfa30):

@ChillOut yes please

OpenStudy (jazzyfa30):

I'm just gonna show you what u got for an answer $1.75x+$1.50y=$525

OpenStudy (chillout):



OpenStudy (chillout):

Just put y in terms of x

OpenStudy (chillout):

and set it to 0, for example. That'd give you x

OpenStudy (jazzyfa30):

350y=0

OpenStudy (chillout):

No, "If I sell 0 candies, how many beverages will I have to sell?"

OpenStudy (jazzyfa30):

300y=0

OpenStudy (chillout):

1.75x + 0*1.5 = 525. < That's what you'd have.

OpenStudy (chillout):

You plug in the value for the amount of candies [y is the amount so it's 0 * the price of candies]

OpenStudy (jazzyfa30):

1.75x=525 because 0*1.50=0

OpenStudy (chillout):

Similarly you can do it for beverages. "If I sell 0 beverages, how many candies will I hve to sell? > 0*1.75 + 1.5y = 525

OpenStudy (chillout):

Correct, @jazzyfa30

OpenStudy (jazzyfa30):

350x+0*1.75=525 1.50x=525 because 0*1.75=0

OpenStudy (chillout):

If you take a look at the graph I drew, It has the solution for 0 candies and 0 beverages. @jazzyfa30, where did that 350x came from? The graph I drew are the solutions for x and y :P

OpenStudy (jazzyfa30):

I did the amount I would have to sell to get 525

OpenStudy (chillout):



OpenStudy (chillout):

Indeed it is right... 1.5y = 525 note, y is the amount, so y = 350 1.75x = 525, so x = 300

OpenStudy (jazzyfa30):

so I did do it right

OpenStudy (chillout):

Can you understand what I'm saying? English is not my 1st language. You mistyped this equation "350x+0*1.75=525" What you want to multiply by "x" is the price, 1.75.

OpenStudy (chillout):

https://www.wolframalpha.com/input/?i=1.75x%2B1.5y%3D525 here is the graph

OpenStudy (jazzyfa30):

that didn't help any it just confused me more

OpenStudy (chillout):

Just ignore the rest. Look at the graph only.

OpenStudy (jazzyfa30):

@johnweldon1993 and @drummmlers

OpenStudy (chillout):

This sums up what we've done. When it intercepts the y axis, we have set x = 0 and when it intercepts the x axis, we have set y = 0

OpenStudy (jazzyfa30):

okay then what is my answer because I'm still not seeing it

OpenStudy (chillout):

My first comment. 1.75x + 1.5y = 525

OpenStudy (jazzyfa30):

I got 3 wrong